A sled of mass 1095 kg is propelled by a rocket. The force exerted by the rocket when it is burning is 7.84×103 N. Assume that the force of friction opposing the motion is 1380 N, what is the acceleration of the sled?
Answer #1

net force = ma

F-f = ma

here F = 7.84*10^10^3 N

f = 1380 N

m = 1095 kg

a = 7840-1380/1095 = 5.89 m/s^2